# Reinforcement Learning for Aligning Large Language Models Agents with Interactive Environments: Quantifying and Mitigating Prompt Overfitting

### v2

@@ Abstract
Mohamed Salim Aissi1, Clement Romac2,3, Thomas Carta3, Sylvain Lamprier4, Pierre-Yves Oudeyer3, Olivier Sigaud1, Laure Soulier1 Nicolas Thome1 1Sorbonne Université, ISIR (MLIA), Paris, France 2Hugging Face 3Inria (Flowers), University of Bordeaux, France 4Univ Angers, LERIA, Angers, France
@@ 1 Introduction
In this paper [^1] Code can be found here: https://anonymous.4open.science/r/Quantifying-and-Mitigating-Prompt-Overfitting-6548, we study the sensitivity of LLMs to prompt formulations and its impact on knowledge acquisition. We define a set of different prompt formulations to evaluate LLM performance when prompts are changed, and also analyze how the LLM represents these prompts. Our findings reveal that the performance of LLMs is highly sensitive to prompt variations, suggesting that fine-tuning only induces superficial updates and fails to improve the acquisition of new knowledge about the environment.  By analogy with observational overfitting in RL (Song et al., 2019), we refer to this phenomenon as prompt overfitting. With this in mind, our paper proposes two main contributions:
@@ 3.1 Problem statement
Figure 2: The fine-tuning framework: we use an LLM as an agent policy in a textual environment $Env$. The $Env$ provides a fixed goal description $g$ for the current episode, a description of the agent’s observation $o$, and a scalar reward $R(o,g)$ for the current step. The goal and observation are formatted using a prompt formulation $P_{i}$. Our experiments reveal that an LLM fine-tuned on $P_{i}$ succeeds with prompts formatted with $P_{i}$ but fails with prompts formatted with $P_{j}$neq $P_{i}$.
On top of the above framework, we study the impact of prompt formulations $P$ =$=\{P_{i}\}_{i\in\{1,\cdots,n\}}$, where $P_{i}:O\times G\mapsto\cal P$$\subset V^{N}$ formats text entries from observations and goals as prompt inputs. We assume that all formulations from $P$ preserve information, i.e. any optimal policy $\pi_{i}^{*}$ using the prompt formulation $P_{i}$ from $P$ can obtain the same amount of rewards as an optimal policy $\pi^{*}$ acting on original observations and goals:  $\forall g\in G,\forall i\in\{1,\cdots,n\},\mathbb{E}_{\tau\sim\pi_{i}^{*}(\tau|g)}[R_{g}(\tau)]=\mathbb{E}_{\tau\sim\pi^{*}(\tau|g)}[R_{g}(\tau)]$.
@@ 3.2 Fine-tuning LLM agents with RL
First, we define a policy $\pi$, based on an LLM for solving tasks in $M$, given any prompt formulation $P_{i}$in $P$. For any goal $g\in{G}$ and observation $o\in O$, we note $\mathbb{P}_{LLM}(w_{k}|$$p_{i}^{o,g}$ $,w_{<k})$ the probability of token $w_{k}$ for prompt $p_{i}^{o,g}$ generated by the prompt formulation $P_{i}$ and the decoding history $w_{<k}$. We also define $\mathbb{P}_{LLM}(a|$$p_{i}^{o,g}$ $)=\Pi_{k=0}^{|a|}\mathbb{P}_{LLM}(w_{k}|$$p_{i}^{o,g}$ $,w_{<k})$ the decoding probability of the sequence corresponding to action $a~{}\in~{}A$ given prompt $p_{i}^{o,g}$. Following (Tan et al., 2024), we use a normalized decoding probability $\mathbb{P}^{norm}_{LLM}(a|$$p_{i}^{o,g}$ $)=\mathbb{P}_{LLM}(a|$$p_{i}^{o,g}$ $)^{\frac{1}{|a|}}$ to better balance actions of various sizes. From these quantities, we build the policy as $\pi(a|$$p_{i}^{o,g}$ $)=\mathbb{P}^{norm}_{LLM}(a|$$p_{i}^{o,g}$ $)/{Z_{i}^{o,g}}$, where $Z^{o,g}_{i}=\sum_{a\in A}\mathbb{P}^{norm}_{LLM}(a|$$p_{i}^{o,g}$ )))) is the partition function. Then, we follow the recent textual RL works for sequential decision-making of Carta et al. (2023) and Tan et al. (2024) to train our LLM agents using Proximal Policy Optimization (ppo(Schulman et al., 2017). Given a subset of prompt formulations $P_{ppo}\subseteq P$, we note $\pi_{P_{ppo}}$ an LLM agent of parameters $\theta$ trained by minimizing the ppo loss $PPO_{loss}(\theta,P_{ppo})$ from rollouts using prompt formulations $P_{ppo}$. Each rollout $\tau$ is obtained for a given prompt formulation $P_{i}$ $\in P_{ppo}$ uniformly sampled from $P_{ppo}$, and is used by the agent during training.
@@ 3.3 Mitigating Prompt Overfitting with Contrastive Learning
$C^{(i,j)}(\theta)=\operatorname*{\mathbb{E}}\limits_{\begin{subarray}{c}(o,g)\sim d^{\pi_{P_{ppo}}}\\ (o^{\prime},g^{\prime})\sim d^{\pi_{P_{ppo}}}\end{subarray}}\max\left(\Delta(z_{\theta}(p^{o,g}_{i}),z_{\theta}(p^{o,g}_{j}))\right.\\ \left.-\Delta(z_{\theta}(p^{o,g}_{i}),z_{\theta}(p^{o^{\prime},g^{\prime}}_{i}))+1),0\right)$ (1)
where $z_{\theta}(p^{o,g}_{i})$ is the latent representation of textual prompt $p_{i}^{o,g}$ produced by prompt formulation $P_{i}$ applied to the observation-goal pair $(o,g)$$\Delta(z,z^{\prime})=||(z)-(z^{\prime})||_{2}^{2}$ is the Euclidean distance between two latent representations $z$ and $z^{\prime}$, and $d^{\pi_{P_{ppo}}}$ is the joint stationary observation-goal distribution of a policy using $P_{ppo}$. In practice, we sample pairs $(o,g)$ from the ppo rollouts, supposed to follow $d^{\pi_{P_{ppo}}}$.
@@
$\vspace{-0.2cm}L(\theta,P_{ppo},P_{c})$ $=$ $PPO_{loss}(\theta,P_{ppo})$
$+$ $\frac{\alpha}{|P_{c}|^{2}}\sum\limits_{P_{i},P_{j}\in P_{c}^{2}}C^{(i,j)}(\theta)\vspace{0.2cm}$
where $\alpha$ is a parameter regulating the impact of $C^{(i,j)}$, $P_{ppo}$ represents the set of prompts used for fine-tuning the LLM policy with the PPO loss, and $P_{c}$ represents the set of prompts used for the contrastive regularization $C^{(i,j)}$. Following (Ni et al., 2022)$C^{(i,j)}$ is only applied to the representation of the first token of the prompt. Appendix C provides implementation details about contrastive regularization.
@@ 4.4 Metrics
$\vspace{-0.2cm}Intra(P_{i})=\frac{1}{|\Gamma|^{2}-|\Gamma|}\sum_{\begin{subarray}{c}(o,g)\in\Gamma\\ \begin{subarray}{c}(o^{\prime},g^{\prime})\in\Gamma\setminus\{o,g\}\end{subarray}\end{subarray}}cos\left(z_{i}^{o,g},z_{i}^{o^{\prime},g^{\prime}}\right).$
@@
$\bullet$ Relevance of Parts of Prompt Information: We study the relationship between inputs and predictions using gradients attribution methods to generate saliency scores for each element of the input (Madsen et al., 2021). For each scenario $\sigma$ and each prompt formulation $P_{i}$, we compute the averaged saliency of prompt tokens for every observation-goal pair $(o,g)$ from $\Gamma$. Saliency scores are computed using the Inseq toolkit (Sarti et al., 2023), as the Integrated Gradient of the output probability of the policy $\pi(a^{*}|$$p_{i}^{o,g}$ )))) with respect to the tokens of the prompt $p_{i}^{o,g}$, with $a^{*}$ the most likely action regarding the observation-goal $o,g$. To enhance interpretability, saliency scores for each prompt tokens are finally aggregated depending on the part of the prompt they belong to, among {possible actions, goal, observation, inventory}. We do so after filtering out the top 5% most significant tokens from each section to eliminate perturbations caused by irrelevant tokens.
@@ 5.2 Q2: State representation
Table 1: Inter and intra-similarity comparison for Flan T5 780M on TWC-Medium on our scenarios. $Intra$ is the average of $Intra($$P_{i}$ )))) over all prompts, $Inter$ is the average of $Inter($$P_{i}$ ,$P_{j}$ )))) over all prompt pairs. For all models, we observe that $Intra$ approaches 1, while $Inter$ is consistently below 0.5. This indicates that the LLMs tend to cluster prompts based on their formulation rather than on content. A similar trend is observed for both $\sigma_{0:3}$ and $\sigma_{0}$ scenarios. Results on other models can be found in Appendix B.2.
@@ 6 Mitigating Overfitting with Contrastive Learning Regularization
Table 3:  Impact on knowledge of the LLM about the environment on TWC QA and TWC OC datasets. * and *** correspond to the p-value (resp. $<0.05$ and $<0.001$) of Welch’s t-test to compare the performance between $\sigma^{{0:3}}_{0}$ and other scenarios. We observe a significant improvement with $\sigma^{{0:3}}_{0}$ scanerio compared to $\sigma_{zs}$, $\sigma_{0}$, and $\sigma_{0:3}$ scenarios across both datasets.
@@ Acknowledgments
This work was supported by the European Commission’s Horizon Europe Framework Programme under grant No 101070381 (PILLAR-robots). Experiments presented in this paper were carried out using the HPC resources of IDRIS under the allocation 2023-[A0151013011] made by GENCI.
@@ B.1 Q1: Prompt sensitivity
[…]
@@
[…]
@@ B.2 Q2: State representation
We also performed a more detailed evaluation of inter and intra distance for models GPT-Neo 1.3B and Flan T5 78M. Table 8 summarizes the results. We can see the same trend for both Flan T5 78M and GPT-Neo 1.3B.
Models
$\sigma_{zs}$
 $\sigma_{0}$  $\sigma_{0:3}$  $\sigma^{{0:3}}_{0}$
$78M$  $Intra$
0.992
 $\pm$ 0.003
0.991
 $\pm$ 0.003
0.991
 $\pm$ 0.003
0.907
 $\pm$ 0.017
 $Inter$
0.376
 $\pm$ 0.019
0.382
 $\pm$ 0.020
0.371
 $\pm$ 0.020
0.806
 $\pm$ 0.029
$780M$  $Intra$
0.998
 $\pm$ 0.001
0.997
 $\pm$ 0.001
0.998
 $\pm$ 0.001
0.939
 $\pm$ 0.017
 $Inter$
0.469
 $\pm$ 0.462
0.458
 $\pm$ 0.449
0.47
 $\pm$ 0.461
0.812
 $\pm$ 0.06
$1.3B$ $Intra$
0.995
 $\pm$ 0.001
0.994
 $\pm$ 0.001
0.997
 $\pm$ 0.001
0.994
 $\pm$ 0.017
$Inter$
0.552
 $\pm$ 0.03
0.539
 $\pm$ 0.01
0.501
 $\pm$ 0.05
0.94
 $\pm$ 0.009
Table 8: Inter and intra-similarity comparison for Flan T5 78M, 780M and GPT-Neo 1.3B models on TWC-Medium on our different scenarios. For all models, we observe that the intra-similarity distance approaches 1, while the inter-similarity distance is consistently below 0.55. This indicates that LLMs tend to cluster prompts based on their formulation rather than their content. A similar trend is observed for both the $\sigma_{0:3}$ and $\sigma_{0}$ scenarios. $\sigma_{0}$ is the agent fine-tuned only on $P_{0}$$\sigma_{0:3}$ is the agent fine-tuned on all prompts and $\sigma^{{0:3}}_{0}$ denotes the agent fine-tuned with the Contrastive loss.

### v3

@@ Abstract
Mohamed Salim Aissi1, Clement Romac2,3, Thomas Carta3, Sylvain Lamprier4, Pierre-Yves Oudeyer3, Olivier Sigaud1, Laure Soulier 1, Nicolas Thome 1,5 1Sorbonne Université, CNRS, ISIR, F-75005 Paris, France 2Hugging Face 3Inria (Flowers), University of Bordeaux, France 4Univ Angers, LERIA, Angers, France 5Institut universitaire de France (IUF) {salim.aissi, laure.soulier, olivier.sigaud, nicolas.thome}@isir.upmc.fr {clement.romac, thomas.carta, pierre-yves.oudeyer}@inria.fr {sylvain.lamprier}@univ-angers.fr
@@ 1 Introduction
In this paper [^1] Code can be found here, we study the sensitivity of LLMs to prompt formulations and its impact on knowledge acquisition. We define a set of different prompt formulations to evaluate LLM performance when prompts are changed, and also analyze how the LLM represents these prompts. Our findings reveal that the performance of LLMs is highly sensitive to prompt variations, suggesting that fine-tuning only induces superficial updates and fails to improve the acquisition of new knowledge about the environment.  By analogy with observational overfitting in RL (Song et al., 2019), we refer to this phenomenon as prompt overfitting. With this in mind, our paper proposes two main contributions:
@@ 3.1 Problem statement
Figure 2: The fine-tuning framework: we use an LLM as an agent policy in a textual environment EnvEnv. The EnvEnv provides a fixed goal description $g$ for the current episode, a description of the agent’s observation $o$, and a scalar reward $R(o,g)$ for the current step. The goal and observation are formatted using a prompt formulation $P_{i}$. Our experiments reveal that an LLM fine-tuned on $P_{i}$ succeeds with prompts formatted with $P_{i}$ but fails with prompts formatted with $P_{j}$ $\neq$ $P_{i}$.
On top of the above framework, we study the impact of prompt formulations $P$ $\{P_{i}\}_{i\in\{1,\cdots,n\}}$, where $P_{i}:O\times G\mapsto\cal P$$\subset V^{N}$ formats text entries from observations and goals as prompt inputs. We assume that all formulations from $P$ preserve information, i.e. any optimal policy $\pi_{i}^{*}$ using the prompt formulation $P_{i}$ from $P$ can obtain the same amount of rewards as an optimal policy $\pi^{*}$ acting on original observations and goals:  $\forall g\in G,\forall i\in\{1,\cdots,n\},\mathbb{E}_{\tau\sim\pi_{i}^{*}(\tau|g)}[R_{g}(\tau)]=\mathbb{E}_{\tau\sim\pi^{*}(\tau|g)}[R_{g}(\tau)]$.
Based on this, we analyze the sensitivity of the policy $\pi$ to variations in prompt formulation compared to the one used during training. We define prompt overfitting as a scenario where, for a given policy $\pi$ trained using a specific formulation $P_{i}$, there exists a prompt formulation $_{j} P\in P$ such that the expected reward for $\pi$ on $P_{i}$ is significantly higher than the expected reward for $\pi$ on $P_{j}$.
@@ 3.2 Fine-tuning LLM agents with RL
First, we define a policy $\pi$, based on an LLM for solving tasks in $M$, given any prompt formulation $_{i} \in PP$. For any goal $g\in{G}$ and observation $o\in O$, we note $\mathbb{P}_{LLM}(w_{k}|$$p_{i}^{o,g}$ $w_{<k})$ the probability of token $w_{k}$ for prompt $p_{i}^{o,g}$ generated by the prompt formulation $P_{i}$ and the decoding history $w_{<k}$. We also define $\mathbb{P}_{LLM}(a|$$p_{i}^{o,g}$ $)=\Pi_{k=0}^{|a|}\mathbb{P}_{LLM}(w_{k}|$$p_{i}^{o,g}$ $w_{<k})$ the decoding probability of the sequence corresponding to action $a~\in~A$ given prompt $p_{i}^{o,g}$. Following (Tan et al., 2024), we use a normalized decoding probability $\mathbb{P}^{norm}_{LLM}(a|$$p_{i}^{o,g}$ $)=\mathbb{P}_{LLM}(a|$$p_{i}^{o,g}$ $)^{\frac{1}{|a|}}$ to better balance actions of various sizes. From these quantities, we build the policy as $\pi(a|$$p_{i}^{o,g}$ $)=\mathbb{P}^{norm}_{LLM}(a|$$p_{i}^{o,g}$ $)/{Z_{i}^{o,g}}$, where $Z^{o,g}_{i}=\sum_{a\in A}\mathbb{P}^{norm}_{LLM}(a|$$p_{i}^{o,g}$ )) is the partition function. Then, we follow the recent textual RL works for sequential decision-making of Carta et al. (2023) and Tan et al. (2024) to train our LLM agents using Proximal Policy Optimization (ppo(Schulman et al., 2017). Given a subset of prompt formulations $P_{ppo}\subseteq P$, we note $\pi_{P_{ppo}}$ an LLM agent of parameters $\theta$ trained by minimizing the ppo loss $PPO_{loss}(\theta,P_{ppo})$ from rollouts using prompt formulations $P_{ppo}$. Each rollout $\tau$ is obtained for a given prompt formulation $P_{i}$ $\in P_{ppo}$ uniformly sampled from $P_{ppo}$, and is used by the agent during training.
@@ 3.3 Mitigating Prompt Overfitting with Contrastive Learning
$C^{(i,j)}(\theta)=\hskip-14.22636pt\operatorname*{\mathbb{E}}\limits_{\begin{subarray}{c}(o,g)\sim d^{\pi_{P_{ppo}}}\\ (o^{\prime},g^{\prime})\sim d^{\pi_{P_{ppo}}}\end{subarray}}\hskip-14.22636pt\max\left(\Delta(z_{\theta}(p^{o,g}_{i}),z_{\theta}(p^{o,g}_{j}))\right.\\ \left.-\Delta(z_{\theta}(p^{o,g}_{i}),z_{\theta}(p^{o^{\prime},g^{\prime}}_{i}))+1),0\right)$ (1)
[…]
@@
$\vskip-5.69046ptL(\theta,P_{ppo},P_{c})\hskip-2.84544pt$ $=$ $PPO_{loss}(\theta,P_{ppo})$
$+$ $\frac{\alpha}{|P_{c}|^{2}}\hskip-2.84544pt\sum\limits_{P_{i},P_{j}\in P_{c}^{2}}\hskip-2.84544ptC^{(i,j)}(\theta)\vskip 5.69046pt$
[…]
@@ 4.4 Metrics
$\vskip-5.69046ptIntra(P_{i})=\frac{1}{|\Gamma|^{2}-|\Gamma|}\hskip-14.22636pt\sum_{\begin{subarray}{c}(o,g)\in\Gamma\\ \begin{subarray}{c}(o^{\prime},g^{\prime})\in\Gamma\setminus\{o,g\}\end{subarray}\end{subarray}}\hskip-14.22636ptcos\left(z_{i}^{o,g},z_{i}^{o^{\prime},g^{\prime}}\right).$
@@
$\bullet$ Relevance of Parts of Prompt Information: We study the relationship between inputs and predictions using gradients attribution methods to generate saliency scores for each element of the input (Madsen et al., 2021). For each scenario $\sigma$ and each prompt formulation $P_{i}$, we compute the averaged saliency of prompt tokens for every observation-goal pair $(o,g)$ from $\Gamma$. Saliency scores are computed using the Inseq toolkit (Sarti et al., 2023), as the Integrated Gradient of the output probability of the policy $\pi(a^{*}|$$p_{i}^{o,g}$ )) with respect to the tokens of the prompt $p_{i}^{o,g}$, with $a^{*}$ the most likely action regarding the observation-goal $o,g$. To enhance interpretability, saliency scores for each prompt tokens are finally aggregated depending on the part of the prompt they belong to, among {possible actions, goal, observation, inventory}. We do so after filtering out the top 5% most significant tokens from each section to eliminate perturbations caused by irrelevant tokens.
@@ 5.2 Q2: State representation
$1.3B$ IntraIntra
0.995
 $\pm$ 0.001
0.994
 $\pm$ 0.001
0.997
 $\pm$ 0.001
0.994
 $\pm$ 0.017
InterInter
0.552
 $\pm$ 0.03
0.539
 $\pm$ 0.01
0.501
 $\pm$ 0.05
0.94
 $\pm$ 0.009
Table 1: Inter and intra-similarity comparison for Flan T5 78M, 780M and GPT-Neo 1.3B models on TWC-Medium on our different scenarios. For all models, we observe that IntraIntra approaches 1, while InterInter is consistently below 0.5. This indicates that the LLMs tend to cluster prompts based on their formulation rather than on content. A similar trend is observed for both $\sigma_{0:3}$ and $\sigma_{0}$ scenarios.
@@ 6 Mitigating Overfitting with Contrastive Learning Regularization
Table 3:  Impact on knowledge of the LLM about the environment on TWC QA and TWC OC datasets. * and *** correspond to the p-value (resp. <0.05<0.05 and <0.001<0.001) of Welch’s t-test to compare the performance between $\sigma^{{0:3}}_{0}$ and other scenarios. We observe a significant improvement with $\sigma^{{0:3}}_{0}$ scanerio compared to $\sigma_{zs}$, $\sigma_{0}$, and $\sigma_{0:3}$ scenarios across both datasets.
@@ Acknowledgments
Experiments presented in this paper were carried out using the HPC resources of IDRIS under the allocation 2024-[A0151013011] made by GENCI. This work was supported by the European Commission’s Horizon Europe Framework Programme under grant No 101070381 (PILLAR-robots) and by PEPR Sharp (ANR-23-PEIA-0008, ANR, FRANCE 2030).
@@ B.1 Q1: Prompt sensitivity
First, we perform the same success rates analysis for LLMs trained with different prompt formulations in two environments. Figure 6 displays these results. Prompt overfitting is also evident when fine-tuning with $P_{1}$$P_{2}$, and $P_{3}$, further supporting our findings. Additionally, we note that the 78M model struggles with BabyAI-Text but shows excellent performance when trained on $P_{3}$, indicating that the LLM can better adapt to certain prompt formulations.
@@
Figure 12: UMAP visualization of the hidden representations of Flan T5 78M with $\sigma_{zs}$ and fine-tuning with  $P_{0}$$P_{1}$$P_{2}$ and $P_{3}$ on BabyAI-Text.